OpenStudy (anonymous):

Find an equation in x and y that has the same graph as the polar equation: r = tan theta

OpenStudy (anonymous):

\[r = \tan \theta\]\[r = \frac{ r \sin \theta }{ r \cos \theta }\]\[r =\frac{ y }{ x }\]But what next?

OpenStudy (psymon):

Can't square both sides?

OpenStudy (anonymous):

\[x^2 + y^2 = \frac{ y^2 }{ x^2 }\] how could you simplify that?

OpenStudy (psymon):

You don't. I'm not going to say I'm certain that's correct, but when you get things like x^2 + y^2 = x, you just leave it, lol.

OpenStudy (anonymous):

Is there a way to put it in terms of y?

OpenStudy (psymon):

Does your question explicitly require it? An answer like that I'd probably just leave as is. I'll be corrected if I'm wrong, but if we have that there is no real simplifying it.

OpenStudy (anonymous):

Well, is\[x^2 + y^2 = \frac{ y^2 }{ x^2 }\]a different way of saying\[y^2 = \frac{ x^4 }{1-x^2 }\]because thats the answer in the back of the book

OpenStudy (psymon):

Ah. Gotcha. Yes, it can be written in that way :3

OpenStudy (anonymous):

Can you help me get it to that point? Please?

OpenStudy (psymon):

\[x ^{2}(x ^{2}+y ^{2})=y ^{2}\] I multiplied x^2 out of the denominator. Now you need to distribute the x^2 and collect all the y terms on one side of the equation.

OpenStudy (psymon):

Find a way to get it?

OpenStudy (anonymous):

Yeah, so \[x^4 + y^2x^2 = y^2\]\[x^4=y^2-y^2x^2\]\[x^4=(1-x^2)y^2\]\[y^2=\frac{ x^4 }{ 1-x^2 }\]

OpenStudy (psymon):

Right ^_^ Just was making sure. Good job.

OpenStudy (anonymous):

Thank you!

OpenStudy (psymon):

mhm :3